\begin{thebibliography}{}
\bibitem{kilbas}
 A.Kilbas and H.Srivatsava and J.Trujillo,
 Theory and Applications of Fractional Differential Equations,
 North Holland
 NY,
 2006.
\bibitem{juma}
G.Jumarie,
Cauchy's integral formula via modified riemann-Liouville derivative for analytic functions of fractional order,
 Appl.Math.Letter,
 23,
2010,
 1444 - 1450.
 
\bibitem{eslami},
M.Eslami,
Exact traveling wave solutions to the fractional coupled nonlinear Schrodinger equations ,
Appl.Math.Comput.,
285,
 2016,
141 - 148.

\bibitem{podlubny}
 I. Podlubny,
 Fractional Differential Equations,
 Academic Press,
 NY,
 1999.

\bibitem{huan},
Huang Q and Zhdanov R.,
 Symmetries and exact solutions of the time fractional Harry-Dym equation with Riemann-Liouville derivative,
Phys. A.,
 409,
2014,
110 - 118.

\bibitem{Bekir}
 A. Bekir,
 New exact traveling wave solutions of some complex nonlinear equations,
 Commun. Nonlinear Sci.,
 14,
 2019,
 1069–1077.
\bibitem{Mirzazadeh}
 M. Mirzazadeh and  M. E. Slami and E. Zerrad et al.,
 Optical solitons in nonlinear directional couplers by sine-cosine function method and Bernoulli’s equation approach,
 Nonlinear Dynam.,
 81,
 2015,
 1933–1949.

\bibitem{Ilhan}
 O. A. Ilhan and A. Esen and H. Bulut  et al,
 Singular solitons in the pseudo-parabolic model arising in nonlinear surface waves,
  Results Phys.,
 12,
 2019,
 1712–1715.
 \bibitem {Alzaidy}
 J. F. Alzaidy,
Fractional Sub-Equation Method and its Applications to the Space-Time Fractional Differential Equations in Mathematical Physics,
Br. J. of Maths. Comp. Sci.,
 2,
 2013,
152-163.
\bibitem{Guo},
S. Guo and Y. Mei and Y. Li and Y. Sun,
The improved fractional sub-equation method and its applications to the space-time fractional differential equations in fluid mechanics,
Phys. Letters A.,
 376,
 2012,
407-411.



 \bibitem{Naher}
 H. Naher and F. A. Abdullah,
 New traveling wave solutions by the extended generalized Riccati equation mapping method of the (2+1)-dimensional evolution equation,
 J. Appl. Math.,
 2012,
 2012,
 486458.
 

 \bibitem{AkbarNaher}
 H. Naher and F. A. Abdullah and M. A. Akbar,
 Generalized and improved $(G'/G)$-expansion method for (3+1) dimensional modified KdV-Zakharov-Kuznetsev equation,
 PLOS One,
 8,
 2013 ,
 64618.
 
 bibitem{Gepreelhomotopy}
  K. A. Gepreel,
 The homotopy perturbation method applied to the nonlinear Kolmogorov-Petrovskii-Piskunov equations,
 Appl.Math.Lett,
 24(8),
 2009,
 1428–1434.

\bibitem{Mahmud}
 F. Mahmud and M. Samsuzzoha and M. A. Akbar,
 The generalized kudryashov method to obtain exact traveling wave solutions of the PHI-four equation and the Fisher equation,
 Results Phys.,
 7,
 2017,
 4296–4302.

\bibitem{bibi}
  S. Bibi and N. Ahmed and U. Khan et al,
 Some new exact solitary wave solutions of the van der Walls model arising in nature,
 Results Phys.,
 9,
 2018,
  648–655,

\bibitem{kaplan}
  M. Kaplan and A. Bekir and A. Akbulut,
  A generalized kudryashov method to some nonlinear evolution equations in mathematical physics,
 Nonlinear Dynam.,
 85,
 2016,
 2843–2850,
 \bibitem{Demiray}
  S. T. Demiray and Y. Pandir and H. Bulut,
 Generalized Kudryashov method for Time-fractional differential equations,
 Abstr. Appl. Anal.,
 2014,
 2014,
 901540

 \bibitem{asmar}
N. H.Asmar,
  Partial Differential Equations with Fourier
Series and Boundary Value Problems, second edition,
 China Machine Press,
Beijing,
2005.
\bibitem{Gepreel}
   K. A. Gepreel and T. A. Nofal and A. A. Alasmari,
 Exact solutions for nonlinear integro-partial differential equations using the generalized Kudryashov method,
 Journal of the Egyptian Mathematical Society,
 25,
 2017,
 438–444.

 \bibitem{kkr}
 K.M. Tamizhmani, K. Krishnakumar and P.G.L.Leach,
 Symmetries and reductions of order for certain nonlinear third and second order differential equations with arbitrary nonlinearity,
 Journal of Mathematical Physics,
56,
2015,
113503-1 - 113503-11.
\bibitem{koparankaplan}
  M. Koparan and M. Kaplan and A. Bekir et al.,
  A novel generalized Kudryashov method for exact solutions of nonlinear evolution equations,
 AIP Conference Proceedings,
  1798,
 2017,
 020082.

\bibitem{bulutpandir}
 H. Bulut and   Y.Pandir and S. T.Demiray.,
  Exact Solutions of Time-Fractional KdV Equations by Using Generalized Kudryashov Method,
 Int. J. Model. Opt.,
 4,
 2014,
 315-320.

\bibitem{bulutbaskonus}
 H. Bulut and  H. M.Baskonus,  and Y.Pandir,
  The modified trial equation method for fractional wave equation and time fractional generalized Burgers equation,
 Abst. Applied Analy.,
 2013,
 1-8.

\bibitem{bulutbernoulli}
 H. Bulut and G. Yel and H. M.Baskonus,
  An application of Improved Bernoulli Sub-Equation method to the nonlinear time fractional generalized Burgers equation,
 Turk.J.Math.Comput.Sci,
  5,
 2015,
 1-7.

\bibitem{elsayed}
 A. M. A. El-Sayed and M. Gaber,
 The Adomian decomposition method for solving partial differential equations of fractal order in finite domains,
 Phys. Lett. A,
  359(3),
 2006,
 175-182.

\bibitem{Gepreelhomotopy}
  K. A. Gepreel,
 The homotopy perturbation method applied to the nonlinear Kolmogorov-Petrovskii-Piskunov equations,
 Appl.Math.Lett,
 24(8),
 2009,
 1428–1434.

\bibitem{Lu}
  B. Lu,
 The first integral method for some time fractional differential equations,
 J.Math.Anal.Appl.,
 395(2),
 2012,
 684-693.

\bibitem{bekirunsal}
  A. Bekir and O. Guner and O.Unsal,
 The first integral method for exact solutions of nonlinear fractional differential equations,
 J.Comp.Nonlinear Dyn,
 10(2),
 2014,
 5 pages.

\bibitem{kkr1}
 K. M. Tamizhmani and K. Krishnakumar and P. G L. Leach,
 Algebraic resolution of equations of the Black–Scholes type with arbitrary time-dependent parameters,
 Applied Mathematics and Computation,
 247,
 2014,
 115-124.
\bibitem{kkr2}
 A. Paliathanasis and K. Krishnakumar and K. M.  Tamizhmani and  P. G. L.  Leach,
 Algebraic resolution of equations of the Black–Scholes type with arbitrary time-dependent parameters,
 Mathematics,
 4(28),
 2016,
 1-14.



\bibitem{Wazwaz}
A. Wazwaz,
The tanh method: Exact solutions of the sine-Gordon and the sinh-Gordon equations,
Appl Math Comput.,
167,
2005,
210-1196.
 
\bibitem{He}
J. H. He and  M. A. Abdou,
New periodic solutions for nonlinear evolution equations using Exp-function method,
Chaos Soliton. Fract.,
34,
2007,
1421–1429.
 

\bibitem {mayelihosseini}
K. Hosseini and P. Mayeli and D. Kumar,
 New exact solutions of the coupled Sine-Gordon equations in nonlinear optics using the modified Kudryashov method,
 J. Mod. Optic.,
65,
2018,
 361–364.
 

\bibitem{mayelihosseiniansari}
K. Hosseini and P. Mayeli and R. Ansari,
Modified Kudryashov method for solving the conformable Time-Fractional Klein-Gordon equations with Quadratic and cubic nonlinearities,
 Optik,
 130,
 2017,
737–742.

\bibitem{bejarbanehhosseini}
K. Hosseini and E. Y. Bejarbaneh and A. Bekir et al,
New exact solutions of some nonlinear evolution equations of pseudo parabolic type,
Opt. Quant. Electron.,
49,
2017,
 241.

\bibitem{egemisirli1}
 S.M.Ege and  E. Misirli,
The modified Kudryashov method for solving some fractional-order nonlinear equations,
 Advances in Difference Equations,
 135,
 2014,
 1-13,
 

\bibitem{egemisirli2}
 S.M.Ege and  E. Misirli,
Solutions of the space-time fractional foam-drainage equation and the fractional Klein-Gordon equation by use of modified Kudryashov method,
 Int. J. of Research Adv. Tech.,
 2321,
 2014,
 384-388.
 
 
\bibitem{ege}
 S.M.Ege and  E. Misirli,
On semianalytical solutions of some nonlinear physical evolution equations with polynomial type auxilary equation,
 PhD Thesis,
 Ege University,
 2015.
 
 
 \bibitem{ryabov}
 P. N. Ryabov and  D. I.Sinelshchikov and M. B.  Kochanov,
 Application of the Kudryashov method for finding exact solutions of the high order nonlinear evolution equations,
 Applied Mathematics and Computation,
 218,
 1999,
 3965-3971,
 
 


 \bibitem{zayed}
 E. M. E. Zayed and K. A. E. Alurrfi,
The modified Kudryashov method for solving some seventh order nonlinear PDEs in mathematical physics,
 World Journal of Modelling and Simulation,
 11,
 2015,
 308-319,
 

 \bibitem{liufu}
 S. Liu and Z. Fu and S. Liu and Q. Zaho,
Jacobi elliptic function expansion method and periodic wave solutions of nonlinear wave equations,
 Phys. Lett. A,
  289,
 2001,
 69-74.
 


 \bibitem{yan}
  Z.Yan,
Abundant families of Jacobi elliptic function solutions of the (2 + 1)-dimensional integrable Davey-Stewartson-type equation via a new method,
 Chaos Solitons Fractals,
 18,
 2003,
 299-309.
 
 
 \bibitem{kudrya}
 N.A. Kudryashov,
]One method for finding exact solutions of nonlinear differential equations,
 Commun.Nonlinear Sci.Number.Simul.,
 17,
 2012,
 2248 - 2253.
 
  \bibitem{pradeep}
 R. G. Pradeep and V. K. Chandrasekar and M. Senthilvelan and M. Lakshmanan,
 On certain new integrable second order nonlinear
differential equations and their connection with two dimensional Lotka-Volterra system,
 J. Math. Phys.,
 51,
 2010,
 033519.

 
 \bibitem{mohanasuba}
  R. Mohanasubha,V. K. Chandrasekar, M. Senthilvelan, and M. Lakshmanan,
 Interplay of symmetries, null forms, Darboux
polynomials, integrating factors and Jacobi multipliers in integrable second-order differential equations,
 Proc. R. Soc. A,
 470(2163),
 2014,
 20130656.
 
 
 \bibitem{Tian1}
 S. F. Tian and Y.F. Zhang and B. L. Feng and H. Q. Zhang,
On the Lie Algebras, Generalized Symmetries and Darboux Transformations of the Fifth-Order Evolution Equations in Shallow Water,
 Chin. Ann. Math. B,
 36,
 2015,
 543-560.
 
 \bibitem{Tian2}
 S. F. Tian and T. T. Zhang and  P. L. Ma and X. Y. Zhang,
 Lie symmetries and nonlocally related systems of the continuous and discrete dispersive long waves system by geometric approach,
 J. Nonlinear Math. Phys.,
 22,
 2015,
 180-193.


 \bibitem{Tian3}
 S. F. Tian and H. Q. Zhang,
On the Integrability of a Generalized Variable-Coefficient
Forced Korteweg-de Vries Equation in Fluids,
 Stud. Appl. Math.,
 132,
 2014,
 214-249.
 
 \bibitem{Galaktionov}
 V. A.Galaktionov  and S.R. Svirshchevskii,
Exact Solutions and Invariant Subspaces of Nonlinear Partial Differential Equations in Mechanics and Physics,
 Chapman and Hall/CRC,
 Florida
 2006

 \bibitem{beck}
 C. Beck and F. Schloegl,
 Thermodynamics of Chaotic Systems,
 Cambridge University Press,
 Cambridge,
 1993.

 \bibitem{Qiao}
 Z. J. Qiao and J. B. Li, 
Negative-order KdV equation with both solitons and kink
wave solutions,
 EPL,
 94,
 2011,
 50003.
 
 
 
 \bibitem{Tang}
 X. Y. Tang and S. Y. Lou and Y. Zhang,
Localized excitations in (2+1)-dimensional systems,
 Phys. Rev. E,
 66,
 2002,
 046601.
 
 \bibitem{korkmazhosseini}
 A. Korkmaz and  K. Hosseini,
Exact solutions of a nonlinear conformable Time-Fractional parabolic equation with Exponential nonlinearity using reliable methods,
 Opt. Quant. Electron.,
 49,
 2017,
 278.
 
 
 
 
 
 
 
 
 
 























% etc
\end{thebibliography}

\RequirePackage{fix-cm}
\documentclass[smallextended]{svjour3}       
\smartqed  
\usepackage{graphicx}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{pdfpages}
\usepackage[none]{hyphenat}
\begin{document}

\title{Exact solutions of time fractional generalized  Burgers-Fisher Equation using generalized Kudryashov method
}


\author{Ramya Selvaraj$^1$        \and
       V. Swaminathan$^1$ \and A.Durga Devi$^2$ \and K.Krishnakumar$^1$ 
}


\institute{${}^1$ 
             Department of Mathematics,
             Srinivasa Ramanujan Centre, 
              SASTRA Deemed to be University, Kumbakonam 612 001, India.
             \email{rsramyaselvaraj@gmail.com, mvsnew@gmail.com}            \\
       ${}^2$
            Department of Physics,
            Srinivasa Ramanujan Centre, 
             SASTRA Deemed to be University, Kumbakonam 612 001, India.
}
\date{Received: date / Accepted: date}

\maketitle

\begin{abstract}
In this article,  we study the generalised Kudryashov method  for the time fractional generalized Burgers-Fisher equation (GBF). Using traveling wave transformation, the time fractional GBF is transformed to nonlinear ordinary differential equation (ODE). Later, in the nonlinear ODE of time fractional GBF,  the generalized Kudryashov and power series method is applied to get exact solutions. 

\keywords{Time fractional differential equation \and Nonlinear differential equation \and Generalized Burgers-Fisher equation \and  Kudryashov method   \and Power series}
\end{abstract}

\section{Introduction}
\hspace{\parindent} The fractional differential equation (FDE) plays a vital role in many branches of science and engineering \cite{kilbas,juma,eslami,podlubny,huan}. FDE has many applications in the field of magnetism, sound waves propagation in rigid porous materials, cardiac tissue electrode interface, theory of viscoelasticity,  fluid mechanics,  lateral and longitudinal control of autonomous vehicles, ultrasonic wave propagation in human cancellous bone,  wave propagation in viscoelastic horn, heat transfer, RLC electric circuit  and so on.
    
    In recent years, for solving time fractional differential equations,
    many researchers have proposed powerful techniques to get an exact solution such as, the sine-cosine method \cite{Bekir,Mirzazadeh}, $G'/G$ expansion method \cite{Ilhan}, the Exp-function method \cite{He}, the tanh method \cite{Wazwaz}, the sub equation method \cite{Alzaidy,Guo}, the improved $G'/G$ expansion method \cite{AkbarNaher}, the invariant subspace method, the generalized Riccati equation method \cite{Naher}, the modified Kudryashov method \cite{bibi,Demiray,Gepreel,bejarbanehhosseini,mayelihosseiniansari,mayelihosseini,kaplan,koparankaplan,korkmazhosseini,Mahmud}  and so on.
   
   Kudryashov method was introduced by Kudryashov \cite{kudrya} for reliable treatment of nonlinear wave equations. For both integer and fractional order, this method is widely used by many researchers such as \cite{bulutbaskonus,ryabov,egemisirli1,egemisirli2,bulutpandir,zayed}. In this work, we apply Kudryashov method to the time fractional generalized Burgers-Fisher equation.
   
   A nonlinear equation which is the combination of reaction, convection and diffusion mechanism is called Burgers-Fisher equation. In the nonlinear equation, the properties of convective phenomenon from Burgers and  diffusion transport as well as reaction kind of characteristics from Fisher are used.  The generalized Burgers-Fisher (GBF) equation  is used in the field of fluid dynamics. It has also been found in some applications such as gas dynamics, heat conduction, elasticity and so on.
   
The aim of this work is organized as follows: In section 2, the algorithmic procedure of generalized Kudryashov method (GKM) is proposed. In section 3, we apply the GKM to find the exact solutions of time fractional generalized Burgers-Fisher equation. Power series method is applied to find the explicit solution of the time fractional GBF equation in section 4 and  section 5 ends with conclusion.

\section{Algorithm of the generalized Kudryashov Method}
 \hspace{\parindent} Let us consider a function $u$ of two real variables, space $x$ and time $t$, then the NLPDE  with fractional order can be written in the form of
\begin{equation} \label{FPDE}
  P(u,D^\alpha_tu,u_x,u_{xx},\ldots)=0
\end{equation}
We illustrate the main steps of generalized Kudryashov method as follows:\\
    In the first step, we obtain the traveling wave solution of Eq.(\ref{FPDE}) of the form:
    \begin{equation}\label{travelingwave}
    u(x,t)=u(\xi),    \;   \xi=kx-\frac{\lambda t^\alpha}{\Gamma[1+\alpha]}, 
\end{equation}
where $k$ and $\lambda$ are arbitrary constants. As a result of this, we obtain a nonlinear ODE in the following form:
\begin{equation} \label{conversionODE}
    N(u,u^\prime,u^{\prime \prime},u^{\prime \prime \prime}, \ldots)=0,
\end{equation}
where the prime indicates differentiation with respect to $\xi$.\\
In the second step, the exact solutions of the nonlinear ODE can be written in the following form:
\begin{equation} \label{divisionodpandq}
u(\xi)=\frac{\sum^K_{i=0}p_iR^i(\xi)}{\sum^N_{j=0}q_jR^j(\xi)}=\frac{P[R(\xi)]}{Q[R(\xi)]},
\end{equation}
where $R$ is $\frac{1}{(1\pm e^\xi)}$. Then 
\begin{equation}
    R_{\xi}=R^2-R.
\end{equation}
Then we obtain
\begin{eqnarray}
u^\prime(\xi)&=&\frac{P^\prime R^\prime Q-PQ^\prime R^\prime}{Q^2}\nonumber\\ 
&=&R^\prime \left[\frac{P^\prime Q-PQ^\prime}{Q^2}\right] \nonumber\\
   &=&(R^2-R)\left[\frac{P^\prime Q-PQ^\prime}{Q^2}\right],\\
 u^{\prime\prime}(\xi)&=&\displaystyle \left[\frac{(R^2-R)}{Q^2}\right]\Bigg[(2R-1)(P^\prime Q-PQ^\prime)\nonumber \\ && +\left(\frac{R^2-R}{Q}\right) [Q(P^{\prime\prime}Q-PQ^{\prime \prime})-2Q^\prime P^\prime P+2P(Q^\prime)^2]\Bigg],\\
 u^{\prime\prime\prime}(\xi)&=&(R^2-R)^3 \Bigg[((P^{\prime\prime \prime}Q-PQ^{\prime\prime \prime}-3P^{\prime\prime}Q^\prime -3Q^{\prime\prime}P^\prime)Q \nonumber \\
 && +6Q(PQ^{\prime\prime}+Q^\prime P^\prime))(Q^3)^{-1}-\frac{6P(Q^\prime)^3}{Q^4}\Bigg] \nonumber \\
 &&+3(R^2-R)^2 (2R-1) \Bigg[\frac{Q(P^{\prime\prime}Q-PQ^{\prime \prime})-2Q^\prime P^\prime P+2P(Q^\prime)^2}{Q^3}\Bigg]
 \nonumber \\
 && +(R^2-R)(6R^2-6R+1)\Bigg[\frac{P^\prime Q-PQ^\prime}{Q^2}\Bigg]
\end{eqnarray}
and so on.\\
  In the third step, we can explain the nonlinear ODE in the form:
   \begin{equation} \label{kudryexactsolution}
       u(\xi)=\frac{p_0+p_1 R+p_2 R^2+\ldots+p_K R^K+\ldots}{q_0+q_1 R+q_2 R^2+\ldots+q_N R^N+\ldots}
   \end{equation}
   By balancing the highest order nonlinear terms  and the highest order derivatives of $u(\xi)$ in Eq.(\ref{conversionODE}) we get some values of $N$ and $K$.\\
    In the fourth step, to calculate the constant $p_i$ and $q_j$, substituting Eq.(\ref{divisionodpandq}) into Eq.(\ref{conversionODE}), it provides a polynomial $R(\xi)$. Establishing the coefficients to zero provides a system of algebraic equations. Solving the system of equations, we can find out the constant. By this way, the exact solution of Eq.(\ref{FPDE}) can be found.

   \section{Applications}
   In this section, we apply the generalized Kudryashov method to the time fractional generalized Burgers-Fisher equation.
    \subsection{Time fractional generalized Burgers-Fisher Equation}
    Let us consider the time fractional generalized Burgers-Fisher equation 
\begin{equation} \label{GBF}
u^\alpha _t+\beta u^\delta u_x-u_{xx}=\gamma u(1-u^\delta) 
\end{equation}
where $0<\alpha \leq 1$, $\alpha$ is the order of the fractional time derivative and $\beta,\gamma,\delta$ are arbitrary constants.
    Now substituting Eq.(\ref{travelingwave}) in Eq.(\ref{GBF}), we obtain
    \begin{equation} \label{tvwreductioneqn}
        k^2 u^{\prime\prime}+(\lambda-k\beta u^{\delta})u^{\prime}+\gamma u(1-u^{\delta})=0
    \end{equation}
Applying the folding transformation
    \begin{equation}
        u(\xi)=v^{\frac{1}{\delta}}(\xi),
    \end{equation}
    we obtain the following equation which is similar to  the  most general form of second order nonlinear oscillator equation \cite{pradeep,mohanasuba} with many arbitrary parameters. With some restrictions on the parameters, they found new integrable equations and further it was discussed by \cite{kkr}. 
\begin{equation}\label{nonlinearode}
   k^2\delta vv^{\prime \prime}+ k^2(1-\delta){v^\prime}^2+( \lambda -k \beta v)\delta v {v^\prime} +\gamma \delta^2 (1-v)v^2=0.
\end{equation}
Balancing the highest power of nonlinear terms of $v^{\prime \prime}$ and $v^3$, then
\begin{equation}
    2K-2N+2=3K-3M \Rightarrow K=N+2.
\end{equation}
Let us choose $N=1,K=3$, then
\begin{equation} \label{riccatifraction}
    v(\xi)=\frac{p_0+p_1R+p_2 R^2+p_3 R^3}{q_0+q_1R},
\end{equation}
\begin{eqnarray}
v^\prime(\xi)&=&(R^2-R) \nonumber \\
&& \times \Bigg[\frac{(p_1+2p_2R+3p_3R^2)(q_0+q_1R)-q_1(p_0+p_1R+p_2 R^2+p_3 R^3)}{(q_0+q_1R)^2}\Bigg], \nonumber\\
v^{\prime \prime}(\xi)&=&\left(\frac{R^2-R}{q_0+q_1 R^2}\right)\Bigg[(2R-1)((p_1+2p_2R+3p_3R^2)(q_0+q_1R^2)\nonumber \\
&&-q_1(p_0+p_1R+p_2 R^2+p_3 R^3))\nonumber \\
&&+\left(\frac{R^2-R}{q_0+q_1R}\right) \Big[(q_0+q_1R)(2p_2+6p_3R)(q_0+q_1R)\nonumber\\
&&-2q_1(p_1+2p_2R+3p_3R^2)(q_0+q_1R) \nonumber \\
&& +2(p_0+p_1R+p_2 R^2+p_3 R^3)q^2_1\Big]\Bigg].
\end{eqnarray}

 The exact solutions of Eq.(\ref{tvwreductioneqn}) are obtained as follows and there are two cases to be considered.
\case 

   For the choices of 

    \[p_2=p_3=0, \;q_0=p_0,\; q_1=0,\; 
    p_1=\frac{k(1+\delta)p_0}{\beta \delta},\] 
    \[\lambda={k^2+k\beta -\gamma \delta}, \;k=\frac{-\beta \delta}{(1+\delta)}\]
and substitutes in Eq.(\ref{riccatifraction}), we get the equation
\begin{eqnarray}
    &&-\beta^2 \delta vv^{\prime \prime}+\beta^2(\delta-1){v^\prime}^2+(\beta^2+\gamma (1+\delta)^2) vv^{\prime}-\gamma(1+\delta)^2 v^3 \nonumber\\
   && -(1+\delta)(\gamma+\gamma \delta+\beta^2 v^\prime)v^2=0.
\end{eqnarray}
The solutions are,
\begin{equation}
 v_1(x,t)=1-\Bigg[\frac{1}{1+ e^{\Big( \frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta)x \Gamma(1+\alpha)}{(1+\delta)^2 \Gamma(1+\alpha)}\Big)}}\Bigg]
\end{equation}
and
\begin{equation}
 v_2(x,t)=1+\Bigg[\frac{1}{1- e^{\Big(\frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta)x \Gamma(1+\alpha)}{(1+\delta)^2 \Gamma(1+\alpha)}\Big)}}\Bigg].
\end{equation}
  Then the exact solutions are,
  \begin{equation} \label{case1exact}
   u_1(x,t)=\Bigg[1-\frac{sec h \Big(\frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta) x \Gamma(1+\alpha)}{2(1+\delta)^2 \Gamma(1+\alpha)} \Big)}{2 e^{\Big(\frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta) x \Gamma(1+\alpha)}{2(1+\delta)^2 \Gamma(1+\alpha)}\Big)}}\Bigg]^\frac{1}{\delta} 
   \end{equation}
   and
   \begin{equation}\label{case1exact1}
   u_2(x,t)=\Bigg[1+\frac{cosec h \Big(\frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta) x \Gamma(1+\alpha)}{2(1+\delta)^2 \Gamma(1+\alpha)} \Big)}{2 e^{\Big(\frac{\beta^2 \delta t^\alpha + \gamma \delta(1+\delta)^2 t^\delta - \beta \delta (1+\delta) x \Gamma(1+\alpha)}{2(1+\delta)^2 \Gamma(1+\alpha)}\Big)}}\Bigg]^\frac{1}{\delta}.
  \end{equation} 
  
 \case 
Considering

    \[p_2=p_3=0,\; q_0=p_0,\;q_1=0,\;
    p_1=\frac{k(1+\delta)p_0}{\beta \delta},\]\\
     \[\lambda={k^2+k\beta -\gamma \delta},\; \gamma=\frac{-k^2(\delta-1)}{\delta^2}\]
and substitutes in Eq.(\ref{riccatifraction}), we get the equation
\begin{equation}
    k vv^{\prime\prime}-k \beta v^\prime v^2+\Big(k+{\beta}\Big)vv^\prime =0.
\end{equation}
The solutions are
\begin{equation}
 v_1(x,t)=1+\frac{2k}{\beta \Big(1+e^{\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{\Gamma(1+\alpha)}}\Big)}
 \end{equation}
 and
 \begin{equation}
 v_2(x,t)=1+\frac{2k}{\beta \Big(1-e^{\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{\Gamma(1+\alpha)}}\Big)}.
\end{equation}
Then the exact solutions are,
\begin{equation} \label{case2exact}
   u_1(x,t)=\displaystyle \Bigg[1+\frac{k sec h \Big(\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{2\Gamma(1+\alpha)}\Big)  }{\beta e^{\Big(\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{2\Gamma(1+\alpha)}\Big) }}\Bigg]^ \frac{1}{\delta} 
   \end{equation}
   and
   \begin{equation} \label{case2exact2}
   u_2(x,t)=\displaystyle\Bigg[1-\frac{k cosec h \Big(\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{2\Gamma(1+\alpha)}\Big)  }{\beta e^{\Big(\frac{k x \Gamma(1+\alpha) - (k^2 + k \beta)t^\alpha}{2\Gamma(1+\alpha)}\Big)} }\Bigg]^ \frac{1}{\delta},
\end{equation}
where $\delta=1$.

\section{Exact Power Series Solutions}
Based on the power series method \cite{Galaktionov,beck,asmar} and symbolic computations \cite{Qiao,Tang,Tian2,Tian1,Tian3}, we  construct the exact power series solutions of Eq.(\ref{GBF}) which is differentiable. First we make use of a transformation
\begin{equation}\label{travwave}
    u(x,t)=u(\eta),       \eta=k x-\frac{\lambda t^\alpha}{\Gamma[1+\alpha]}, 
\end{equation}
where $k$ and $\lambda$ are arbitrary constants with k, $\lambda \neq 0$. Substituting Eq.(\ref{travwave}) into Eq.(\ref{GBF}), then we can get the nonlinear ODE 
\begin{equation}\label{nonlinearode1}
    k^2\delta vv^{\prime \prime}+k^2(1-\delta){v^\prime}^2+( \lambda -k \beta v)\delta v {v^\prime} +\gamma \delta^2 (1-v)v^2=0.
\end{equation}
We suppose that Eq.(\ref{nonlinearode}) has the following solution:
\begin{equation}\label{newpowerseries}
    v(\xi)=\sum^\infty_{n=0} q_n \xi^n,
\end{equation}
where $q_n (n=0,1,2,...)$ are constants. Then, we have 
\begin{eqnarray}\label{newpowerseries1}
v^{\prime} (\xi)=\sum^\infty_{n=0}(n+1) q_{n+1} \xi^n,
\end{eqnarray}
\begin{eqnarray}\label{newpowerseries2}
v^{\prime \prime} (\xi)=\sum^\infty_{n=0}(n+1)(n+2) q_{n+2} \xi^n.
\end{eqnarray}
Substituting Eq.(\ref{newpowerseries}),Eq.(\ref{newpowerseries1}) and Eq.(\ref{newpowerseries2}) into Eq.(\ref{nonlinearode1}) and solving, we get 
\begin{equation}\label{qtwo}
    q_2=\frac{1}{2}\Big[\Big(\frac{\delta-1}{\delta}\Big)\frac{q_1^2}{q_0}-\frac{\lambda q_1}{k^2}+\frac{\beta q_0 q_1}{k}-\frac{\gamma \delta q_0}{k^2}+\frac{\gamma \delta q_0^2}{k^2}\Big], 
\end{equation}
when $n=0$.
When $n\geq1$, we obtain
\begin{eqnarray}\label{qn}
q_{n+2}&=&\frac{1}{(n+2)(n+1)}\Big[\Big(\frac{\delta-1}{\delta}\Big)\frac{(n+1){q^2_{n+1}}}{q_n}-\frac{\lambda (n+1)q_{n+1}}{k^2}\nonumber\\
&&+\frac{\beta (n+1)q_n q_{n+1}}{k}-\frac{\gamma \delta q_n}{k^2}+\frac{\gamma \delta q_n^2}{k^2}\Big].
\end{eqnarray}

It is easy to prove the convergence of the power series
Eq.(\ref{newpowerseries}) with the coefficients given in Eq.(\ref{qtwo}) and Eq.(\ref{qn}). Therefore this power series solution of Eq.(\ref{newpowerseries}) is an exact analytic solution.\\
 Hence, the power series solution of Eq.(\ref{newpowerseries}) can be written as follows:

\begin{eqnarray}
v(\xi)&=&q_0+q_1(\xi)+q_2(\xi^2)+\sum^\infty_{n=1}q_{n+2}\xi^{n+2}
\nonumber \\
&=&q_0+q_1(\xi)+\frac{1}{2}\Big[\Big(\frac{\delta-1}{\delta}\Big)\frac{q_1^2}{q_0}-\frac{\lambda q_1}{k^2}+\frac{\beta q_0 q_1}{k}-\frac{\gamma \delta q_0}{k^2}+\frac{\gamma \delta q_0^2}{k^2}\Big]\xi^2   \nonumber\\
&&+\sum^\infty_{n=1}\frac{1}{(n+2)(n+1)}\Big[\Big(\frac{\delta-1}{\delta}\Big)\frac{(n+1){q^2_{n+1}}}{q_n}-\frac{\lambda (n+1)q_{n+1}}{k^2}\nonumber\\&&+\frac{\beta (n+1)q_n q_{n+1}}{k}-\frac{\gamma \delta q_n}{k^2}+\frac{\gamma \delta q_n^2}{k^2}\Big]\xi^{n+2}.
\end{eqnarray}

\section{Conclusion}
In this article,  the time fractional generalized Burgers-Fisher equation have been transformed into a nonlinear ordinary differential equation by using folding transformation. The resultant equation  is similar to the most familiar general form of second order nonlinear oscillator equation  with some restrictions on the parameters. Then we applied generalized Kudryashov method and also power series method to the resultant equation for finding  exact solutions. Plots are given for the exact solutions which show the dynamics of solutions with suitable parametric choices. 

\begin{acknowledgements}
The authors thank the Department of Science and Technology-Fund Improvement of S\&T Infrastructure in Universities and Higher Educational Institutions Government of India (SR/FST/MSI-107/2015) for carrying out this research work.
\end{acknowledgements}




\end{document}